OpenStudy (anonymous):

y=3x-1;{(2,5),(-2,7),(0,-1),(1,1)}

OpenStudy (anonymous):

just substitute the x coordinate into the x part and solve for y. Start with the first one, (2,5). 3*2-1 = y = 5. so that one is right. 3*-2-1 = -7 so that's not right. 3*0-1 = -1 so that's right. and 3*1-1 = 2 that's not right. so i don't know if you are supposed to get two right answers but there are.